OpenStudy (anonymous):

one pipe can fill a tank in 20 minutes, while another takes 30 minutes to fill the same tank. How long would it take the two pipes together to fill the tank?

OpenStudy (dumbcow):

look at rates per minute pipe 1: rate of 1/20 per minute pipe 2: rate of 1/30 per minute both pipes: 1/20 + 1/30 per minute --> 1/20 + 1/30 = 3/60 + 2/60 = 5/60 = 1/12 --> 12 minutes